Please help me, I can manage regular problems but I'm not very good at word problems. I can't seem to make heads or tails of this one.
An object suspended from an ideal spring is pulled away from it’s equilibrium position and is released, causing the object to osculate up and down. The Frequency (f, in s^-1) of oscillation is related to the mass (m in kg) of the object, and to K a constant that depends on the stiffness of the spring, by the equation
f = the square root of k/m
If the frequency is 5.000 s^-1 and the spring constance is 16. 00 kg/s^, what is the mass of the object?

Answer by scott8148(6628)

5=sqrt(16/m) ___ squaring gives 25=16/m ___ 25m=16 ___ m=16/25 ___ m=.64 kg